### v1

# Measuring Sample Importance in Data Pruning for Training LLMs from a Data Compression Perspective

Abstract
Compute-efficient training of large language models (LLMs) has become an important research problem. In this work, we consider data pruning as a method of data-efficient training of LLMs, where we take a data compression view on data pruning. We argue that the amount of information of a sample, or the achievable compression on its description length, represents its sample importance. The key idea is that, less informative samples are likely to contain redundant information, and thus should be pruned first. We leverage log-likelihood function of trained models as a surrogate to measure information content of samples. Experiments reveal a surprising insight that information-based pruning can enhance the generalization capability of the model, improves upon language modeling and downstream tasks as compared to the model trained on the entire dataset.
1 Introduction & Related Work
Training large language models (LLM) requires high computational costs due to their immense sizes and commensurately large datasets Kaplan et al. [2020], Hoffmann et al. [2022]. Recently, computationally efficient methods to train deep models based on data pruning have gained interest. Data pruning concerns deciding which samples are important for training, and remove unimportant samples from training datasets. Sorscher et al. [2022] proposed that neural scaling law Kaplan et al. [2020] may be overcome with data pruning. Their method uses distances between samples and their cluster centroids in the embedding space to measure the sample importance. Tan et al. [2024] proposed to assess the sample importance by efficiently measuring the change in empirical risk when the sample is removed from the training set. The aforementioned works apply data pruning to labeled data for supervised learning or focus on visual understanding tasks. By contrast, data pruning methods for LLMs have been largely underexplored. Instead, text deduplication methods have been proposed  Raffel et al. [2020], Abbas et al. [2023] to remove redundant data based on semantic similarity. However, they do not quantify the importance of samples for data pruning.
Meanwhile, connections between LLMs and data compression have been recently explored. Ge et al. [2023] proposed In-Context Autoencoder for context compression. Deletang et al. [2023] extensively study the view of LLM as a model for data compression. The predictive power of LLMs can be used for an optimally efficient expression of data based on the prediction probability, so as to compress various types of multi-modal data.
In this paper, we propose to leverage the compression capability of LLMs for data pruning. If a model learns the distribution of corpus, it can estimate the amount of information of a sample, i.e., the optimal number of bits required to compress the sample. Importantly, if the amount of information of a sample is low, the sample likely contains information which is redundant or appear frequently in dataset. Less informative samples are regarded as less important and thus are removed from dataset. To estimate the information of samples, we first train called data probe model which is a small model trained with a subset of corpus. We use the log-likelihood output of data probe model as a surrogate to measure the sample information. Next, we prune the dataset based on the estimates by data probe model, and train a target model with the pruned dataset. Experimental results are quite surprising: in many cases, the proposed pruning actually improves the model performance. In some cases, the performance of generation and/or downstream tasks are maintained up to pruning 50% of the pretraining corpus. The key insight is that, the proposed pruning based on sample information helps removing irrelevant data, which promotes the generalization capability of language models.
2 Method
2.1 Compression View to Data Pruning
We take a data compression view to data pruning for language models as follows. Let $p(\cdot)$ denote the true distribution of words from a corpus. Given a sample $W=(w_{1}w_{2}\ldots w_{n})$ which is a sequence of words $\{w_{i}\}$, consider
$H(W,p)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{1}{p(w_{i}|w_{<i})}$
We view $H(W,p)$ as related to the minimum description length required to compress words in $W$. Indeed, the optimal compression based on entropy, e.g, Huffman coding Huffman [1952] or arithmetic coding Rissanen [1976], uses approximately $\lceil\log_{2}1/p(w_{i}|w_{<i})\rceil$ bits to encode $w_{i}$ given previous words. Alternatively, $H(W,p)$ represents the amount of information per word in passage $W$. Consider a next-word predictor $q$, i.e., a language model. The negative log-likelihood of $W$ on model $q$, which is our main metric for sample importance, is given by
$H(W,q)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{1}{q(w_{i}|w_{<i})}$ (1)
From an information theoretic view, (1) represents the description length of words in $W$ using sub-optimal distribution $q$. Meanwhile, if we average $nH(W,q)$ over $W$, it is the log-loss used for training language models. Thus, training is equivalent to minimizing the description length of, i.e., compressing, the corpus data Deletang et al. [2023].
We hypothesize that, if $q$ is well-trained to approximate $p$, we may use $H(W,q)$ as a surrogate for $H(W,p)$. In particular, $H(W,q)$ can provide an upper bound of $H(W,p)$ in an average sense. Indeed, we have
$\mathcal{H}(p)$ $=\mathbf{E}_{W\sim p}\left[H(W,p)\right]$
$\mathcal{H}(p,q)$ $=\mathbf{E}_{W\sim p}\left[H(W,q)\right]$
where $\mathcal{H}(p)$ and $\mathcal{H}(p,q)$ denotes the per-word entropy and cross-entropy with respect to $q$, respectively. The cross-entropy is an upper bound of entropy Cover [1999], i.e.,
$\mathcal{H}(p)\leq\mathcal{H}(p,q)$
That is, we always need extra bits to encode words with distribution $p$ when $q$ is used for encoding.
Thus, if $q$ is sufficiently trained to achieve low log-loss (cross-entropy) for next-word prediction, $H(W,q)$ is likely to provide a good upper bound on $H(W,p)$ for sufficiently long $n$. In turn, we hypothesize that sample $W$ with low $H(W,q)$ (implying low $H(W,p)$) is of low informational value, because it has short description length. That is, $W$ likely contains word/phrases which appear frequently in the corpus, i.e., contains redundant information. Thus, it makes sense to remove samples with small $H(\cdot,q)$ values from the dataset.
Another view is that, one can show that the perplexity of $W$ for model $q$ is given by $e^{H(W,q)}.$ Thus, low $H(W,q)$ results in low perplexity or “surprisal”, where the surprisal of a sample is a term used for the information content of the sample in an information theoretic context.
2.2 Pruning Method
Overview. We first train reference model $q$, called data probe model. The goal is to economically measure $H(\cdot,q)$ of the samples. To that end, we only use a subsample of the pretraining corpus to train $q$. Next, we prune the samples with low $H(\cdot,q)$ from the dataset. The pruned dataset is used for training the target LLM.
Data probe model. We train a data probe model denoted by $q$. To reduce computational costs, we train $q$ on a subsample of the pre-training corpus. Still, the subsample size should be sufficient so that $H(\cdot,q)$ provide a good bound on the description lengths. We heuristically determine the subsample size by hypothesizing that the model is sufficiently trained if the rate of decrease in the training loss saturates, similar to the early stopping method.
As a result, the subsample size is set to be about 12% of training dataset: see Fig. 1. Later experiments show that the proposed level of training is sufficient for a good pruning performance.
Pruning. Next, we perform pruning on the dataset $\mathcal{D}$. We first measure $H(W,q)$ for all $W\in\mathcal{D}$. In our setup, $W$ is a fixed-length sequence. Denote the target pruning ratio by $\eta$. Prune bottom $\eta$of samples in terms of $H(\cdot,q)$ values from $\mathcal{D}$. Use the pruned dataset to train the LLM.
Figure 1: Loss curve of GPT-125M over 1 sweep of the training dataset. For data efficiency, we stopped training the probe model at the point where the decrease in loss saturates, i.e., at about 12% of the entire dataset.
3 Experiment
3.1 Experimental Settings
As pretraining corpus, we use a subset of c4 Raffel et al. [2020] via random sampling, which is a total of about 3 billion tokens from GPT-2 tokenizer Radford et al. [2019]. We use decoder-only transformers Vaswani et al. [2017] using GPT-2 tokenizer as the language model. As the data probe model, we use a 125M-parameter model trained on 0.3 billion tokens randomly sampled from the total corpus. As the target model, we use both 125M- and 345M-parameter models. We evaluate the language modeling of target models on One billion words Chelba et al. [2014] and wikitext-103 Merity et al. [2016] where the models are pre-trained on different datasets. The downstream tasks for the target models are evaluated using glue benchmark Wang et al. [2018]. Detailed hyperparameters are in Appendix A.1.
3.2 Pretraining Loss with Pruning
Fig 2 shows the test loss of the 345M target model during pretraining. Interestingly, pruning 10% achieves a lower and pruning 20% obtains a similar test loss, as compared to no-pruning case. As we will see later, the lower loss indeed leads to improved performances in language modeling and downstream tasks with pruning.
Figure 2: Test loss of 345M GPT per training token.
Pruning % Proposed Random High Entropy 0 90.23 10 89.32 90.95 98.47 20 86.88 93.40 100.06 30 87.09 99.89 109.51 40 84.19 93.73 107.48 50 91.11 97.08 104.39 60 91.24 101.32 103.30 70 102.93 104.65 152.94 80 101.70 110.39 181.16 90 147.18 165.39 284.59
Table 1: Perplexity of 125M per pruning ratio on One Billion Words corpus. Bold denotes the best performance for each pruning ratio, and the underline indicates better performances than no-pruning case.
Pruning % Proposed Random High Entropy 0 52.23 10 50.59 51.56 54.71 20 51.71 52.25 57.99 30 53.46 54.97 61.35 40 55.06 56.91 71.03 50 58.26 59.63 74.50 60 61.97 64.29 85.91 70 62.66 70.22 97.87 80 71.46 85.42 126.69 90 115.03 135.34 212.67
Table 2: Perplexity of 125M per pruning ratio on wikitext-103 corpus. Bold denotes the best performance for each pruning ratio, and the underline indicates better performances than no-pruning case.
3.3 Language Modeling
We evaluate language modeling of target models using two test corpus. Table 1 shows the perplexity of each model per pruning ratio on One Billion Words corpus Chelba et al. [2014]. The proposed pruning significantly outperforms the random pruning. Surprisingly, the proposed pruning achieves better performance up to 40% pruning ratio, and on-par performance up to 60% ratio, as compared to no-pruning case. We observe a similar trend with different corpus: Table 2 shows the perplexity results of language modeling on wikitext-103 corpus Merity et al. [2016]. The results show that, our pruning can actually improve the generalization capability of the target model, which we explain as follows. Our pruning removes samples that are less informative or contain highly redundant information. Such redundancy in the dataset may cause overfitting such samples similar to duplication, i.e., the model may get biased towards such information, hampering generalization. Thus, pruning based on estimates of information content of samples may alleviate overfitting problems. For comparison, we tested removing samples with high $H(\cdot,q)$ first: see “High Entropy” columns in Table 1 and 2. Interestingly, it performs much worse than random pruning, which conversely shows the effectiveness of the proposed importance metric of samples.
3.4 Downstream tasks
Figure 3: Text classification results on CoLA dataset in glue benchmark. Left: 125M, Right: 345M. The green line indicates the performance without pruning.
Figure 4: Textual similarity results on SST2 dataset in glue benchmark. Left: 125M, Right: 345M. The green line indicates the performance without pruning.
We finetune and evaluate each model in downstream tasks such as text classification and textual similarity tasks from glue Wang et al. [2018]. The top of Fig. 3 shows the text classification results of 125M models on CoLA dataset, which compare between the proposed method and random pruning. The results show that, the accuracy of our method is always above that of the random pruning. Next, we scaled up the target model to 345M, and the bottom of Fig. 3 shows that the proposed data pruning method achieves significant performance gains over random pruning. Notably, the proposed pruning outperforms no-pruning case with almost up to 50% pruning ratio, i.e., pruning helps improving the performance of language models.
Fig. 4 illustrates the results of the textual similarity task on SST2 dataset for 125M and 345M models. For both models, while random pruning continually decreases performances, the proposed method achieves similar or even better accuracy than no-pruning case up to 50% pruning ratio.
Overall, we make a similar observation to pretraining case: the proposed pruning not only reduces the training cost of language models, but also improves the performance of downstream tasks. We conclude that pruning based on the information content of samples can improve the generalization capabilities of the models on downstream tasks.
4 Conclusion
In this paper, we took an information-theoretic view of compression for data pruning in training language models. We use the log-likelihood function of the model to estimate the compressed description length of a sample, or its informativeness. Pruning based on the proposed estimate of sample importance enables removing samples with redundant information, which not only reduces computational costs, but also, surprisingly, enhances the generalization capability of language models. Experiments with various corpus and tasks validated the effectiveness of the proposed pruning.
Appendix A Appendix
A.1 Hyperparameters
Hyperparameter Value Model size 125M, 345M Pretrain learning rate 5e-4 Pretrain Learning rate scheduler Cosine Pretrain warmup ratio 0.01 % Pretrain weight decay 0.1 Pretrain batch size 64 Sequence length 1024 Finetuning learning rate 2e-5 Finetuning weight decay 0.01 Finetuning batch size 32 Optimizer Adam Kingma and Ba [2014]
Table 3: Detailed hyperparameters.